RC meeting and Run Plan, Wednesday December 8th

From SBS wiki
Jump to: navigation, search

NEXT RC IS BRIAN :)

RC Meeting Information

Run Coordinator: Eric King

Location: Counting House Conference Room

Time: 3:15PM

Remote connection link: https://bluejeans.com/256205934

Toll-Free Number (U.S.& Canada): 1-888-240-2560

Bluejeans CODE: 256 205 934

Upcoming Shifts Which Need To Be Filled

A couple more shifts were signed up for. A total of 12 shifts remain until the winter break shutdown half of which are this upcoming weekend.

  1. Sat Dec 11th, OWL, SL
  2. Sat Dec 11th, DAY, SL
  3. Sat Dec 11th, DAY, TO
  4. Sat Dec 11th, SWING, TO
  5. Sun Dec 12th, OWL, SL
  6. Sun Dec 12th, DAY, TO
  7. Sun Dec 12th, SWING, TO
  8. Wed Dec 15th, OWL, SL
  9. Fri Dec 17th, OWL, TO
  10. Sat Dec 18th, OWL, TO
  11. Sat Dec 18th, DAY, TO
  12. Sun Dec 19th, DAY, TO

Accelerator and Hall Status

  • (Hall) Access during Tuesday SWING [1] to look at GEMS (timing depended solely on GEM work).
    • GEM experts completed work in attempt to address the low efficiency in UVa X/Y GEM-1 [2] (review results in subsystem updates on Wednesday).
    • Sebastian completed some work on HCal and BBCal [3] (review in subsystem updates on Wednesday).
  • (Hall) Access for Andrew (fr. Jessie's crew) to install the aluminum absorber (perhaps have results to review in meeting).

To Be Done During Extended Down Time

  • (GEM) Would like to take special lower-HV test runs while beam is off.
  • (BBCal) May potentially need more cosmics data.
  • (HCal) Would like to take special data runs.

Run Plan

Continue taking production data. :) We should be entering cycle 10 soon.

  • (Not updated yet) Bogdan will update/extend run plans available on the SBS Wiki.

Data Collected (RC Week)

Data collected by shift.

Week 12/1 thru 12/8
Triggered Events (Millions)
Day Shift LH2 LD2 Dummy
Dec 1 SWING 10.0 - -
Dec 2 OWL 1.0 20.2 -
DAY - 60.6 -
SWING 20.5 5.6 2.7
Dec 3 OWL - 43.4 -
DAY 20.4 27.0 -
SWING - 38.2 -
Dec 4 OWL - 48.2 -
DAY 7.9 40.9 -
SWING - 44.0 3.5
Dec 5 OWL 13.0 40.8 -
DAY - 47.3 -
SWING 12.9 30.1 -
Dec 6 OWL - 50.2 -
DAY 12.1 30.3 -
SWING - 30.1 -
Dec 7 OWL 6.0 20.0 3.5
DAY - 30.0 -
SWING - 18.4 -
Dec 8 OWL 6.0 42.0 -
DAY - 61.4 -
Total Triggered Events: 103.8 728.9 9.7

By Cycle:

Week 12/1 thru 12/8
Triggered Events (Millions)
CYCLE LH2 LD2 Dummy
1 11.0 80.8 2.7
2 20.5 76.0 -
3 20.4 86.4 -
4 7.9 84.9 3.5
5 13.0 88.1 -
6 12.9 80.3 -
7 12.1 80.4 3.5
8 6.0 80.4
9 6.0 71.4
Totals 103.8 728.9 9.7

Opportunistic Access Requests (Running List)

  • [Ralph] Timing hodoscope (remove card which will stay in hall).
  • [Maria/Todd] GRINCH (Could use any reasonable amount of time [LED runs]).
  • Removal of INFN J2 from Hall [Jessie has been contacted about this; will be completed over winter break barring an opportunistic access before then].
  • [Sebastian] Swap out SD card in RaspPi in HCal upper crate. Needs 1+ hour of controlled access.

Unresolved Issues

  • [Resolution underway] Hodoscope HV alarms going off [4]; I created a HALOG entry and tagged Ralph/Rachael [5]. Apparent repeat of issue on 11/25 [6]. Ralph came in to double check the issue and make sure it wasn’t a hodoscope hardware problem. Ralph will be coordinating with Stephen Wood on a resolution.

Subsystem Updates

  1. HCal
    1. Better Panguin plots made.
    2. Golden runs added to plots.
  2. BBCal
    1. Nothing of note.
  3. Timing Hodoscope
    1. Nothing new.
  4. GRINCH
    1. HV settings may be too low (can be handled by Alex).
  5. GEMS
    1. Andrew fixed the low efficiency problem--software related.
    2. J0: Turned on HV this morning... still strange behavior on mid and bottom modules.
    3. Andrew has had a change of opinion on potentially altering the HV by -50V.
  6. Beamline
    1. No new updates.
  7. Replay analysis
    1. Nothing of note.